# FaithSCAN: Model-Driven Single-Pass Hallucination Detection for Faithful Visual Question Answering

### v2

@@ I Introduction
Visual Question Answering (VQA) is a representative image-to-text (I2T) task in which a model generates natural language answers given an image and a corresponding question [19]. Recent advances in multimodal large language models (MLLMs) and vision-language models (VLMs) have enabled impressive integration of visual perception and language understanding, supporting increasingly complex open-ended VQA tasks [12, 43]. Despite impressive performance, they frequently generate fluent yet ungrounded or incorrect responses, a phenomenon known as hallucination [21, 9, 3], which in VQA settings manifests as responses misaligned with the input visual evidence or contradictory to factual knowledge [29, 18]. Among them, faithfulness hallucinations [19], where answers violate visual grounding, are particularly critical in VQA, since reliable factual verification depends on the model first attending to and correctly interpreting the image [27]. Consequently, detecting faithfulness hallucinations is a necessary step for ensuring answer reliability and user trust, motivating the development of effective detection methods that can identify unfaithful answers (exmaples illustrated in Fig. 1).
@@
The remainder of this paper is organized as follows. Section II reviews related work. Section III formally defines the problem setting and introduces the key concepts used throughout this study. Section IV presents the proposed methodology. Section V describes the experimental setup, including datasets, evaluation metrics, and implementation details. Section VI reports experimental results and analysis. Finally, Section VII concludes the paper.
@@ II-B Uncertainty-Based Hallucination Detection
Model-internal uncertainty has emerged as an effective and relatively inexpensive cue for hallucination detection. Both language models and VLMs exhibit increased predictive uncertainty when hallucinations occur [46, 37, 1, 15]. Methods such as semantic entropy (SE) [9] and likelihood-based measures [32, 48, 22] quantify uncertainty by aggregating variations across multiple decoding outputs. In VQA, extensions incorporate visual perturbations or multimodal semantic clustering, e.g., vision-amplified semantic entropy [28] and dense geometric entropy [10], which estimate hallucination likelihood by comparing semantic distributions under clean and perturbed visual inputs. These methods demonstrate that uncertainty signals can effectively capture hallucinations, including in safety-critical domains such as medical VQA [28]. Nevertheless, most approaches rely on repeated sampling or perturbations to obtain reliable estimates, which increases inference cost and can be sensitive to stochastic decoding behavior.
@@ VI-A Overall Effectiveness and Generalization (RQ1)
We evaluate FaithSCAN under in-distribution (ID) and out-of-distribution (OOD) settings across four datasets. FaithSCAN leverages token-level uncertainty signals (log-likelihood, entropy, embeddings) and visual patch features, with sequence-level representations obtained via global average pooling and fused through attention (see Section VI-C for architectural choices).
@@ VI-A1 In-distribution Evaluation
ID results (Table III) show that FaithSCAN reaches or surpasses SOTA performance, achieving up to 8% AUROC improvement and maintaining top performance across metrics and backbone models. Sampling-based baselines such as SEU, SelfCheckGPT and $P(\text{True})$ perform competitively but incur higher sampling and inference costs, whereas FaithSCAN achieves strong detection in a single pass. SE underperforms due to long-form, detailed answers that undermine clustering assumptions. Performance further improves with more capable backbones, suggesting richer internal representations provide clearer uncertainty signals.
@@ VI-A2 Out-of-distribution Evaluation
Figure 3: RQ2: Illustrative examples of the consistency judgment process based on model-driven supervision, showing both positive and negative cases. (a) The sample is drawn from POPE, where the LLM-based judger correctly identifies a contradiction relationship between the model-generated answer and the ground-truth reference. (b) The sample is drawn from VQA v2, where the generated answer is not strictly precise; nevertheless, it is judged as entailment by the LLM judger.
@@ VI-B1 Human Evaluation
TABLE V: RQ2: Consistency analysis between model-driven supervision signals and ground-truth hallucination annotations across different datasets.
@@
HalLoc-VQA 0.960 0.834 0.846
POPE 0.990 0.872 0.879
HaloQuest 0.960 0.816 0.829
VQA v2 0.930 0.796 0.787
Quantitatively (Table V), the supervision achieves high agreement with human labels (0.960–0.990), with Cohen’s Kappa and MCC exceeding 0.81 across datasets. POPE shows the highest consistency due to its simpler object-level hallucinations, while more complex datasets like HalLoc-VQA, HaloQuest, and VQA v2 yield slightly lower but still strong agreement.
@@ VI-B2 Automated Evaluation
Figure 4: Distributions of post-hoc reliability scores on three datasets. Answers are from IB-T5-XL; bars are colored by the original three-way hallucination labels (Sec. IV-C1). Results for other backbones are in Supplementary Material D.2.
@@ VI-C3 Layer-wise Representation Analysis
We evaluated token-level representations from different decoder layers for hallucination detection on HalLoc-VQA using IB-T5-XL, LLaVA-8B, and Qwen3-8B. Each layer was tested on 100 samples, repeated three times, using all uncertainty signals from Section IV-D with attention-based fusion. Fig. 7 shows that representations from the earliest and deepest layers consistently underperform, while intermediate layers achieve higher performance across all models. The layer-wise uncertainty exhibits fluctuations, indicating that different layers encode complementary information: intermediate layers balance token-level variability and semantic abstraction, capturing cross-modal cues relevant to hallucination. This trend is particularly clear for Qwen3-8B and LLaVA-8B.
@@ VII Conclusion
This paper investigates faithfulness hallucination detection in VQA from a model internal perspective. Our results indicate that many hallucinations are already encoded in the internal states of VLMs during generation, rather than emerging solely at the output level. Uncertainty signals from language decoding, visual perception, and cross modal interaction provide complementary views of different hallucination failure pathways, and their relative contributions vary across model architectures and hallucination types.
More broadly, this work suggests that hallucinations in VQA are structured phenomena rooted in multimodal reasoning dynamics. The effectiveness of lightweight and trainable detectors operating on frozen model states demonstrates that reliable hallucination detection can be achieved without repeated sampling or external verification. We hope these findings motivate future research to treat hallucination as an intrinsic property of vision language reasoning, supporting the development of more reliable and interpretable multimodal systems.
@@ B.2 Extraction of Uncertainty Signals
  • token-level hidden representations from (emb) a designated layer of the language model;

### v3

@@ I Introduction
Visual Question Answering (VQA) is a representative image-to-text (I2T) task in which a model generates natural language answers given an image and a corresponding question [19]. Recent advances in multimodal large language models (MLLMs) and vision-language models (VLMs) have enabled impressive integration of visual perception and language understanding, supporting increasingly complex open-ended VQA tasks [12, 43]. Despite impressive performance, they frequently generate fluent yet ungrounded or incorrect responses, a phenomenon known as hallucination [21, 9, 3], which in VQA settings manifests as responses misaligned with the input visual evidence or contradictory to factual knowledge [29, 18]. Among them, faithfulness hallucinations [19], where answers violate visual grounding, are particularly critical in VQA, since reliable factual verification depends on the model first attending to and correctly interpreting the image [27]. Consequently, detecting faithfulness hallucinations is a necessary step for ensuring answer reliability and user trust, motivating the development of effective detection methods that can identify unfaithful answers (examples illustrated in Fig. 1).
@@
The remainder of this paper is organized as follows. Section II reviews related work. Section III defines the problem setting and introduces key concepts. Section IV presents the proposed methodology. Section V describes the experimental setup, including datasets, metrics, and implementation. Section VI reports results and analysis. Finally, Section VII concludes the paper.
@@ II-B Uncertainty-Based Hallucination Detection
Model-internal uncertainty has emerged as an effective and relatively inexpensive cue for hallucination detection. Both language models and VLMs exhibit increased predictive uncertainty when hallucinations occur [46, 37, 1, 15]. Methods such as SE [9] and likelihood-based measures [32, 48, 22] quantify uncertainty by aggregating variations across multiple decoding outputs. In VQA, extensions incorporate visual perturbations or multimodal semantic clustering, e.g., vision-amplified semantic entropy [28] and dense geometric entropy [10], which estimate hallucination likelihood by comparing semantic distributions under clean and perturbed visual inputs. These methods demonstrate that uncertainty signals can effectively capture hallucinations, including in safety-critical domains such as medical VQA [28]. Nevertheless, most approaches rely on repeated sampling or perturbations to obtain reliable estimates, which increases inference cost and can be sensitive to stochastic decoding behavior.
@@ VI-A Overall Effectiveness and Generalization (RQ1)
We evaluate FaithSCAN under in-distribution (ID) and out-of-distribution (OOD) settings across four datasets. FaithSCAN leverages token-level uncertainty signals (log-likelihood, entropy, embeddings) and visual patch features. Sequence-level representations are obtained via global average pooling and fused through attention (see Section VI-C for architectural choices).
@@ VI-A1 In-distribution Evaluation
ID results (Table III) show that FaithSCAN reaches or surpasses SOTA performance, achieving up to 8% AUROC improvement and maintaining top performance across metrics and backbone models. Sampling-based baselines such as SEU, SelfCheckGPT, and $P(\text{True})$ perform competitively but incur higher sampling and inference costs, whereas FaithSCAN achieves strong detection in a single pass. SE underperforms due to long-form, detailed answers that undermine clustering assumptions. Performance further improves with more capable backbones, suggesting richer internal representations provide clearer uncertainty signals.
@@ VI-A2 Out-of-distribution Evaluation
Figure 3: RQ2: Illustrative examples of the consistency judgment process based on model-driven supervision, showing both positive and negative cases. (a) The sample is drawn from POPE, where the LLM-based judger correctly identifies a contradiction relationship between the model-generated answer and the ground-truth reference. (b) The sample is drawn from VQA v2, where the generated answer is not strictly precise; nevertheless, it is judged as entailment by the LLM judge.
@@ VI-B1 Human Evaluation
TABLE V: RQ2: Consistency analysis between model-driven supervision signals and human annotations across different datasets.
@@
HalLoc-VQA 0.960 0.834 0.837
POPE 0.990 0.947 0.948
HaloQuest 0.960 0.858 0.861
VQA v2 0.950 0.798 0.799
Quantitatively (Table V), the proposed supervision exhibits strong agreement with human annotations, achieving agreement scores in the range of 0.950–0.990. Moreover, Cohen’s Kappa and MCC are consistently high across datasets, with most values exceeding 0.8.
@@ VI-B2 Automated Evaluation
Figure 4: RQ2: Distributions of post-hoc reliability scores on three datasets. Answers are from IB-T5-XL; bars are colored by the original three-way hallucination labels (Sec. IV-C1). Results for other backbones are in Supplementary Material Sec. D.2.
@@ VI-C3 Layer-wise Representation Analysis
We evaluate token-level representations from different decoder layers for hallucination detection on HalLoc-VQA using IB-T5-XL, LLaVA-8B, and Qwen3-8B. Each layer is tested on 100 samples, repeated three times, using all uncertainty signals from Section IV-D with attention-based fusion. Fig. 7 shows that representations from the earliest and deepest layers consistently underperform, while intermediate layers achieve higher performance across all models. The layer-wise uncertainty exhibits fluctuations, indicating that different layers encode complementary information: intermediate layers balance token-level variability and semantic abstraction, capturing cross-modal cues relevant to hallucination. This trend is particularly clear for Qwen3-8B and LLaVA-8B.
@@ VII Conclusion
This paper investigates faithfulness hallucination detection in VQA from a model internal perspective. Our results indicate that many hallucinations are already encoded in the internal states of VLMs during generation, rather than emerging solely at the output level. Uncertainty signals from language decoding, visual perception, and cross-modal interaction provide complementary views of different hallucination failure pathways. Their relative contributions vary across model architectures and hallucination types.
More broadly, this work suggests that hallucinations in VQA are structured phenomena rooted in multimodal reasoning dynamics. The effectiveness of lightweight and trainable detectors operating on frozen model states demonstrates that reliable hallucination detection can be achieved without repeated sampling or external verification. We hope these findings motivate future research to treat hallucination as an intrinsic property of vision–language reasoning, supporting the development of more reliable and interpretable multimodal systems.
@@ B.2 Extraction of Uncertainty Signals
  • token-level hidden representations (emb) from a designated layer of the language model;